House price y is estimated as a function of the square footage of a house x and a dummy variable d that equals 1 if the house has ocean views. The estimated house price, measured in $1,000s, is given by yˆy^ = 118.90 + 0.12x + 52.60d. ***NOTE yˆy^=Y-hat

a-1. Compute the predicted price (in $1,000s) of a house with ocean views and square footage of 2,000. (Round intermediate calculations to at least 4 decimal places. Round your answer to 2 decimal places.)

a-2. Compute the predicted price (in $1,000s) of a house with ocean views and square footage of 3,000. (Round intermediate calculations to at least 4 decimal places. Round your answer to 2 decimal places.)

b-1. Compute the predicted price (in $1,000s) of a house without ocean views and square footage of 2,000. (Round intermediate calculations to at least 4 decimal places. Round your answer to 2 decimal places.)

b-2. Compute the predicted price (in $1,000s) of a house without ocean views and square footage of 3,000. (Round intermediate calculations to at least 4 decimal places. Round your answer to 2 decimal places.)

c. Discuss the impact of ocean views on the house price.

  • An ocean view increases the value of house price by approximately 52.60 thousand dollars.

  • An ocean view decreases the value of house price by approximately 10% at the given level.
Answer #1

Solution

y^ = 118.90 + 0.12x + 52.60d

a1) x = 2000 and d = 1

y = 118.90 + 0.12*2000 + 52.60 = 411.50

a2) x = 3000 and d = 1

y = 118.9 + 0.12*3000+52.60 = 531.50

b1) x = 2000 and d = 0

y = 118.90+0.12*2000 = 358.90

b2) x = 3000 and d = 0

y = a2) - 52.60 = 531.5-52.6 = 478.90

c) An ocean view increases the value of house price by approximately 52.60 thousand dollars